# Comparative Performance Analysis of Quantum Machine Learning Architectures for Credit Card Fraud Detection

### v1

@@ Abstract
As financial fraud becomes increasingly complex, effective detection methods are essential. Quantum Machine Learning (QML) introduces certain capabilities that may enhance both accuracy and efficiency in this area. This study examines how different quantum feature map and ansatz configurations affect the performance of three QML-based classifiers—the Variational Quantum Classifier (VQC), the Sampler Quantum Neural Network (SQNN), and the Estimator Quantum Neural Network (EQNN)—when applied to two non-standardized financial fraud datasets. Different quantum feature map and ansatz configurations are evaluated, revealing distinct performance patterns. The VQC consistently demonstrates strong classification results, achieving an F1 score of 0.88, while the SQNN also delivers promising outcomes. In contrast, the EQNN struggles to produce robust results, emphasizing the challenges presented by non-standardized data. These findings highlight the importance of careful model configuration in QML-based financial fraud detection. By showing how specific feature maps and ansatz choices influence predictive success, this work guides researchers and practitioners in refining QML approaches for complex financial applications.
Index Terms:
Quantum Machine Learning, Quantum Neural Networks, Fraud Detection, Quantum Variational Classifier
@@ I Introduction
Credit card fraud represents a pervasive and escalating global threat, costing individuals and organizations billions of dollars each year. For instance, a 2022 UK Finance study reported fraud-related losses of £1.2 billion, with online solicitations accounting for approximately 80% of fraudulent activities [1]. Similarly, a CNBC report published in 2023 estimated that fraud imposed an \$8.8 billion burden on American consumers in 2022 [2], while a joint assessment by the European Banking Authority and the European Central Bank indicated that credit card fraud reached 633 million euros in the first half of 2023 [3]. These escalating figures threaten consumers and financial institutions, contributing to broader economic instability.
@@
Conventional fraud detection methods have advanced significantly, yet they face persistent challenges. The increasing volume and complexity of transactional data, the highly imbalanced nature of fraud detection datasets, and the growing need for real-time decision-making continue to strain classical Machine Learning (ML) approaches [4, 5]. Quantum Machine Learning (QML) offers a promising avenue to overcome these limitations [6], potentially enhancing computational speed, handling complex data distributions more effectively, and improving overall detection accuracy. Moreover, QML has demonstrated success across a wide range of applications—extending from healthcare and finance to climate modeling [7, 8, 9], materials science [10, 11], and cybersecurity—illustrating its versatility in tackling complex [12], data-intensive problems [13, 14]. This success largely depends on selecting and optimizing various QML models and architectural configurations designed to employ the unique capabilities of quantum computing [15, 16, 17]. Consequently, developing effective QML-based solutions necessitates a thorough investigation of different model architectures and configurations to identify those that most effectively address the specific challenges of each application domain. In the context of financial fraud detection, this involves analyzing how different QML architectures impact performance metrics such as accuracy, recall, and F1-score, thereby ensuring the development of robust and efficient detection systems [18].
@@
This study focuses on examining how various QML configurations—specifically, configurations of feature maps and ansatz architectures—influence accuracy, recall, and F1-score when applied to two real-world credit card fraud detection datasets. By evaluating a range of model setups and pinpointing those configurations that achieve the best performance, this work provides valuable insights for designing more effective QML-based fraud detection systems. Ultimately, this analysis contributes to the broader goal of protecting consumers and businesses from increasingly sophisticated financial threats.
@@
  • A comprehensive analysis of three emerging QML models—Variational Quantum Classifier (VQC), Sampler Quantum Neural Network (SQNN), and Estimator Quantum Neural Network (EQNN)—applied to credit card fraud detection tasks.
  • A thorough empirical evaluation involving two credit card fraud detection datasets, assessing the responses and performance of the QML models.
  • An in-depth examination of the impact of different feature maps and ansatz configurations on model accuracy and detection capabilities.
  • Insights into identifying the optimal feature map–ansatz configurations that offer the most effective performance in quantum-enhanced financial fraud detection.
@@
Figure 1: Overview of the proposed methodology. Two benchmark datasets undergo initial preprocessing, after which three QML models are evaluated. Each model is tested with three distinct feature map architectures, and each selected feature map is further assessed using four different ansatz configurations
@@ II Literature Review
Despite significant progress in classical approaches, these methods face persistent challenges due to the volume, complexity, and continuously evolving patterns of fraudulent activities. This has motivated interest in QML, which employs quantum properties—including entanglement and superposition—to operate efficiently in high-dimensional feature spaces. As a result, QML has the potential to process large datasets more rapidly, capture complex interactions among transactional features, and express intricate patterns, offering tangible benefits for fraud detection tasks [26, 27, 28].
@@ III Background
Figure 2: Fundamental architecture of a QML model. It begins with the transformation of classical data into quantum states, represented as $|\psi_{i}\rangle$. These states are processed by an ansatz $\mathbf{U}(\theta)$, which applies quantum gates such as rotation and entanglement. The model incorporates a classical optimizer that iteratively updates the parameters of the ansatz based on a loss function until convergence is achieved. Finally, the output of the ansatz is measured to provide the model’s predictions for various tasks.
@@ III-A3 Z Feature Map
The Z feature map encoding method also offers parameters like entity size, repetition count, and entanglement approach crucial for circuit inspection. Since there are no entanglement gates in the Z feature map, the entanglement strategy is zero. By offering a different quantum feature map that fits particular usage scenarios where entangled gates must be avoided, this feature map enhances the ZZ feature map method.
@@ IV-A Variational Quantum Classifiers
Finally, measurement is performed on the transformed quantum state $\left|\Psi(\theta)\right\rangle$ to extract information and compute a classical output. Steps 2 and 3 are iteratively executed to minimize the loss function $J(\theta)$. The process terminates when an acceptance criterion, such as convergence of the loss function, is met [43].
Figure 5: Architecture of the VQC model, it consists of a feature map, the $U$ operator is then applied in a measurement to obtain the classical output $\hat{y}$. This process is repeated where a classical optimizer is applied to update the parameters $\theta$ until the loss function $J(y;\hat{y})$ is minimized.
@@ IV-B Estimator Quantum Neural Networks
Figure 6: Architecture of the EQNN model. The process starts with a feature map, followed by an ansatz and measurements to extract classical data. This data is then passed through a fully connected classical neural network, which computes the output value $\hat{y}$. A classical optimizer iteratively updates the parameters $\theta^{*}$, $\mathbf{W}^{*}$, and $b^{*}$ to minimize the loss function $J(y,\hat{y})$.
The operation of the EQNN begins with data preparation via a quantum feature map. Classical data, denoted as $\mathbf{x}=(x_{1},x_{2},\dotsc,x_{n})$, is encoded into quantum states $\left|\Psi_{0}(\theta)\right\rangle$ through the quantum feature map $\Phi:\mathbf{x}\mapsto\left|\Psi_{0}(\theta)\right\rangle$. Following the encoding stage, the quantum states $\left|\Phi(\mathbf{x})\right\rangle$ are processed through the ansatz to transform the encoded states into output quantum states $\left|\Psi(\theta)\right\rangle$. Subsequently, measurements are performed on the output states $\left|\Psi(\theta)\right\rangle$, extracting classical features from the quantum system. These features are obtained by measuring the qubits in the computational basis ${\left|0\right\rangle,\left|1\right\rangle}$. The classical features obtained from the measurement stage are then processed by a fully connected classical neural network, which computes the output values $\hat{y}$. This step combines the strengths of classical and quantum processing for robust learning. Finally, the model undergoes an optimization phase where the parameters of both the quantum and classical components are adjusted. The goal is to determine the optimal parameters $\theta^{*}$ for the ansatz and the weights $\mathbf{W}^{*}$ and biases $b^{*}$ of the classical neural network by minimizing the loss function $J(y;\hat{y})$. This process can be formulated as:
$(\theta^{*};\mathbf{W}^{*};b^{*})=\arg\min_{\theta,\mathbf{W},b}J(y;\hat{y}).$ (6)
@@ V-A Dataset and Feature Selection
  • “Age”: Categorizes customer age into bands: ’0’ ($\leq 18$), ’1’ (19–25), ’2’ (26–35), ’3’ (36–45), ’4’ (46–55), ’5’ (56–65), ’6’ ($>65$), and ’U’ (Unknown).
@@
This dataset is unbalanced, containing 492 fraud cases out of 284,807 transactions, with the positive class representing only 0.172% of all transactions. To address the class imbalance in both datasets, “Random Under Sampling” was implemented to create balanced datasets with a 50/50 ratio. Each dataset was reduced to 492 fraudulent and 492 non-fraudulent transactions. PCA was applied to the European dataset, reducing its dimensions to seven features named “V1” through “V7”.
@@ V-C BankSim Dataset
Table II highlights the performance of the VQC model. Notably, the ZZ and Pauli feature maps paired with the Efficient SU2 ansatz yield the highest F1-scores of 0.68 and 0.71, respectively. These results suggest that introducing entanglement through the feature map improves the VQC’s ability to learn complex interactions. Fig. 12 supports this, as most configurations stabilize after oscillatory dynamics, indicating successful optimization. Similarly, Table III reveals the robust performance of the SQNN model. The Pauli feature map combined with Two Local and Real Amplitudes achieves F1-scores of 0.84 and 0.85, respectively, showcasing the effectiveness of these configurations. Fig. 13 demonstrates consistent convergence behavior across these configurations, with some achieving minimal loss values after initial oscillations. In contrast, Table IV shows that the EQNN model generally underperforms compared to VQC and SQNN. While the Pauli feature map with Pauli Two Design achieves a peak F1-score of 0.59, configurations involving the Z feature map yield significantly lower scores, underscoring the importance of entanglement in feature maps. Fig. 17 corroborates these findings, representing varying levels of convergence stability across configurations.
@@
SQNN ZZ Real Amplitudes 0.74 0.77 0.61 0.68
@@
EQNN ZZ Real Amplitudes 0.45 0.46 0.48 0.37
@@ V-D European Dataset
SQNN ZZ Real Amplitudes 0.65 1 0.23 0.37
@@
EQNN ZZ Real Amplitudes 0.45 0.45 0.47 0.40
@@ V-D2 Analysis of Ansatz
V-E Discussion
The analytical results highlight the critical interaction between feature maps, ansatz configurations, and dataset characteristics in shaping model performance. The ability of a model to balance the performance metrics is heavily influenced by the capacity of the feature map to encode meaningful data relationships and the ansatz to exploit these encoded patterns effectively. This interaction varies across datasets and model architectures, emphasizing the importance of a tailored approach to configuration selection. The performance variations across models reflect their inherent architectural differences. VQC consistently leverages expressive ansatz-feature map combinations to achieve high precision and balanced F1-scores, especially on the European dataset. SQNN demonstrates robust adaptability, excelling in configurations that pair computational efficiency with high representational power, such as the Z feature map with Real Amplitudes on the BankSim dataset. In contrast, EQNN’s limitations in handling complex patterns suggest that either its architectural simplicity or constraints in its quantum circuit design impede its ability to capitalize on the available data encodings. These disparities underscore the importance of aligning model architectures with feature map and ansatz choices to maximize performance.
Additionally, the results suggest that while entanglement-enhanced feature maps like ZZ and Pauli can add value, their success is contingent on the model’s ability to process the increased complexity. Models like SQNN benefit significantly from such feature maps when paired with flexible ansatzes like Two Local, which provide a rich search space to navigate complex data relationships. Conversely, simpler feature maps like Z demonstrate versatility and effectiveness, likely due to their ability to reduce computational noise and simplify optimization. This analysis underscores the necessity of a strategic, data-driven approach to selecting feature map-ansatz configurations. The interplay between dataset characteristics and model architecture determines the optimal balance between computational efficiency and performance, guiding the design of more effective quantum and hybrid quantum models for fraud detection.

### v3

@@ Abstract
As financial fraud becomes increasingly complex, effective detection methods are essential. Quantum Machine Learning (QML) introduces certain capabilities that may enhance both accuracy and efficiency in this area. This study examines how different quantum feature maps and ansatz configurations affect the performance of three QML-based classifiers, the Variational Quantum Classifier (VQC), the Sampler Quantum Neural Network (SQNN), and the Estimator Quantum Neural Network (EQNN), when applied to two non-normalized financial fraud datasets. Different quantum feature map and ansatz configurations are evaluated, revealing distinct performance patterns. The VQC consistently demonstrates strong classification results, achieving an F1-score of 0.88, while the SQNN also delivers promising outcomes. In contrast, the EQNN struggles to produce robust results, emphasizing the challenges presented by non-standardized data. Statistical validation using ANOVA confirms the significance of observed performance differences. Additionally, robustness tests on the best-performing models under five quantum noise types show that they maintain competitive performance, supporting their practical applicability. These findings highlight the importance of careful model configuration in QML-based financial fraud detection. By showing how specific feature maps and ansatz choices influence predictive success, this work guides researchers and practitioners in refining QML approaches for complex financial applications.
@@ I Introduction
Credit card fraud represents a pervasive and escalating global threat, costing individuals and organizations billions of dollars each year. For instance, a 2022 UK Finance study reported fraud-related losses of £1.2 billion, with online solicitations accounting for approximately 80% of fraudulent activities [39]. Similarly, a CNBC report published in 2023 estimated that fraud imposed an \$8.8 billion burden on American consumers in 2022 [7], while a joint assessment by the European Banking Authority and the European Central Bank indicated that credit card fraud reached 633 million euros in the first half of 2023 [4]. These escalating figures threaten consumers and financial institutions, contributing to broader economic instability. Machine Learning (ML) has been widely applied across domains for classification and decision tasks [52, 49], including fraud detection. Conventional fraud detection methods have advanced significantly, yet they face persistent challenges. The increasing volume and complexity of transactional data, the highly imbalanced nature of fraud detection datasets, and the growing need for real-time decision-making continue to strain ML approaches [5, 32, 1, 44]. Quantum Machine Learning (QML) offers a promising avenue to overcome these limitations [42], potentially enhancing computational speed, handling complex data distributions more effectively, and improving overall detection accuracy. Moreover, QML has demonstrated success across a wide range of applications, extending from healthcare and finance to climate modeling [12, 23, 17], materials science [9, 20], and cybersecurity, illustrating its versatility in tackling complex [14], data-intensive problems [34, 26]. This success largely depends on selecting and optimizing various QML models and architectural configurations designed to employ the unique capabilities of quantum computing [22, 18, 21]. Consequently, developing effective QML-based solutions necessitates a thorough investigation of different model architectures and configurations to identify those that most effectively address the specific challenges of each application domain. In the context of financial fraud detection, this involves analyzing how different QML architectures impact performance metrics such as accuracy, recall, and F1-score, thereby ensuring the development of robust and efficient detection systems [19].
@@
This study focuses on examining how various QML configurations, specifically, configurations of feature maps and ansatz architectures, influence accuracy, recall, and F1-score when applied to two real-world credit card fraud detection datasets. In addition, we incorporate statistical validation of performance through Analysis of Variance (ANOVA) and evaluate the robustness of the most effective models under different quantum noise scenarios. By evaluating a range of model setups and pinpointing those configurations that achieve the best performance, this work provides valuable insights for designing more effective QML-based fraud detection systems. Our study further provides a systematic multi-factor analysis of QML architectures, exploring how model type, feature map, ansatz design, and dataset characteristics interact and affect performance. This broader evaluation uncovers dependencies that are often overlooked in prior work and clarifies how these architectural components collectively influence fraud detection outcomes. Ultimately, this analysis contributes to the broader goal of protecting consumers and businesses from increasingly sophisticated financial threats.
@@
  • A comprehensive analysis of three emerging QML models, Variational Quantum Classifier (VQC), Sampler Quantum Neural Network (SQNN), and Estimator Quantum Neural Network (EQNN), applied to credit card fraud detection.
  • A thorough empirical evaluation of two real-world credit card fraud datasets, examining how each QML model behaves across key performance metrics.
  • An in-depth examination of how feature maps and ansatz configurations influence model accuracy and detection capability, showing how encoding choices and circuit expressivity affect performance.
  • A rigorous statistical validation using ANOVA to measure the individual and interaction effects of model type, dataset, feature map, and ansatz on F1-score, clarifying which design choices drive performance.
  • Insights into optimal feature map–ansatz combinations that deliver the strongest results for quantum-enhanced fraud detection.
  • A robustness analysis of the best-performing model from each dataset, evaluated under five types of quantum noise to assess resilience and transferability to real NISQ hardware conditions.
@@
Figure 1: Overview of the proposed methodology. Two benchmark datasets undergo initial preprocessing, after which three QML models are evaluated. Each model is tested with three distinct feature map architectures, and each selected feature map is further assessed using four different ansatz configurations.
@@ II Literature Review
Despite significant progress in classical approaches, these methods face persistent challenges due to the volume, complexity, and continuously evolving patterns of fraudulent activities. This has motivated interest in QML, which employs quantum properties—including entanglement and superposition—to operate efficiently in high-dimensional feature spaces. As a result, QML has the potential to process large datasets more rapidly, capture complex interactions among transactional features, and express intricate patterns, offering tangible benefits for fraud detection tasks [36, 45].
@@ III Background
Figure 2: Fundamental architecture of a QML model. It involves transforming classical data into quantum states $|\psi_{i}\rangle$, processing them through an ansatz $\mathbf{U}(\theta)$, and optimizing its parameters using a classical optimizer based on a loss function until convergence is achieved iteratively. The final quantum state is measured to generate predictions.
@@ III-A3 Z Feature Map
The Z feature map encoding method also offers parameters like repetition count, and entanglement approach, crucial for circuit inspection. Since there are no entanglement gates in the Z feature map, the entanglement strategy is zero. By offering a different quantum feature map that fits particular usage scenarios where entangled gates must be avoided, this feature map enhances the ZZ feature map method.
@@ IV-A Variational Quantum Classifiers
Finally, measurement is performed on the transformed quantum state $\left|\Psi(\theta)\right\rangle$ to extract information and compute a classical output. Steps 2 and 3 are iteratively executed to minimize the loss function $J(\theta)$. The process terminates when an acceptance criterion, such as convergence of the loss function, is met.
Figure 5: Architecture of the VQC model, it consists of a feature map, the $U$ operator is then applied in a measurement to obtain the classical output $\hat{y}$. A classical optimizer iteratively updatethe parameters $\theta$ to minimize the loss function J(y;y^)J($\hat{y}$).
@@ IV-B Estimator Quantum Neural Networks
Figure 6: Architecture of the EQNN model. The process starts with a feature map, processing it via ansatz, then feeds the measured output into a classical neural network. A classical optimizer updates both quantum and neural network parameters ($\theta^{*}$, $\mathbf{W}^{*}$, and $b^{*}$) to minimize the loss function J(y,y^)J($\hat{y}$).
The operation of the EQNN begins with data preparation via a quantum feature map. Classical data, denoted as $\mathbf{x}=(x_{1},x_{2},\dotsc,x_{n})$, is encoded into quantum states $\left|\Psi_{0}(\theta)\right\rangle$ through the quantum feature map $\Phi:\mathbf{x}\mapsto\left|\Psi_{0}(\theta)\right\rangle$. Following the encoding stage, the quantum states $\left|\Phi(\mathbf{x})\right\rangle$ are processed through the ansatz to transform the encoded states into output quantum states $\left|\Psi(\theta)\right\rangle$. Subsequently, measurements are performed on the output states $\left|\Psi(\theta)\right\rangle$, extracting classical features from the quantum system. These features are obtained by measuring the qubits in the computational basis $\left|0\right\rangle,\left|1\right\rangle$}. The classical features obtained from the measurement stage are then processed by a fully connected classical neural network, which computes the output values $\hat{y}$. This step combines the strengths of classical and quantum processing for robust learning. Finally, the model undergoes an optimization phase where the parameters of both the quantum and classical components are adjusted. The goal is to determine the optimal parameters $\theta^{*}$ for the ansatz and the weights $\mathbf{W}^{*}$ and biases $b^{*}$ of the classical neural network by minimizing the loss function J(y;y^)J($\hat{y}$). This process can be formulated as:
[…]
@@ V-A Dataset and Feature Selection
  • “Age”: Categorizes customer age into bands: ’0’ ($\leq 18$), ’1’ (19–25), ’2’ (26–35), ’3’ (36–45), ’4’ (46–55), ’5’ (56–65), ’6’ (>65>65), and ’U’ (Unknown).
@@
This dataset is unbalanced, containing 492 fraud cases out of 284,807 transactions, with the positive class representing only 0.172% of all transactions. To address the class imbalance in both datasets, “Random Under Sampling” is implemented to create balanced datasets with a 50/50 ratio. Each dataset is reduced to 492 fraudulent and 492 non-fraudulent transactions. PCA is applied to the European dataset, reducing its dimensions to seven features named “V1” through “V7”.
@@ V-C BankSim Dataset
Table II highlights the performance of the VQC model. Notably, the ZZ and Pauli feature maps paired with the Efficient SU2 ansatz yield the highest F1-scores of 0.68 and 0.71, respectively. These results suggest that introducing entanglement through the feature map improves the VQC’s ability to learn complex interactions. Fig. 12 supports this, as most configurations stabilize after oscillatory dynamics, indicating successful optimization. Similarly, Table III reveals the robust performance of the SQNN model. The Pauli feature map combined with Two Local and Real Amplitudes achieves F1-scores of 0.84 and 0.85, respectively, showcasing the effectiveness of these configurations. While Fig. 13 demonstrates consistent convergence behavior across these configurations, with some achieving minimal loss values after initial oscillations. In contrast, Table IV shows that the EQNN model generally underperforms compared to VQC and SQNN. While the Pauli feature map with Pauli Two Design achieves a peak F1-score of 0.59, configurations involving the Z feature map yield significantly lower scores, underscoring the importance of entanglement in feature maps. Fig. 17 corroborates these findings, representing varying levels of convergence stability across configurations.
@@
SQNN $Z$ Real Amplitudes 0.74 0.77 0.61 0.68
@@
EQNN $Z$ Real Amplitudes 0.45 0.46 0.48 0.37
@@ V-D European Dataset
SQNN $Z$ Real Amplitudes 0.65 1 0.23 0.37
@@
EQNN $Z$ Real Amplitudes 0.45 0.45 0.47 0.40
@@ V-D2 Analysis of Ansatz
V-E Statistical Analysis
To confirm the validity of the differences observed between the QML model architectures, we perform a one-way ANOVA test on the F1-score. ANOVA evaluates whether the means of multiple groups differ significantly by comparing the variance between groups to the variance within groups, expressed as an F-value. A larger F-value and a smaller p-value ($p<0.05$) indicate that the differences are unlikely due to random chance. This analysis considers four main experimental factors: the model, the feature map, the ansatz, and the dataset.
TABLE VIII: Results of the one-way ANOVA test for the F1-score.
Factor F p-value Interpretation
Dataset 1.71 0.1917 Not significant
Model 335.08 $1.18\times 10^{-102}$ Highly significant
Feature Map 8.96 $1.44\times 10^{-4}$ Significant
Ansatz 10.96 $7.51\times 10^{-23}$ highly significant
As shown in Table VIII, the results reveal that the choice of QML model architecture has a statistically significant effect on F1-score performance. Specifically, the “Model” factor shows a very high F-value (335.08) and an extremely low p-value ($1.18\times 10^{-102}$), indicating a highly significant influence. The feature map and ansatz factors also yield significant F-values (8.96 and 10.96, respectively), with corresponding p-values well below 0.05, confirming their impact. In contrast, the Dataset factor is not significant (F = 1.71, $p=0.1917$), suggesting consistent performance across datasets.
To support these results, Fig. 18 presents the distribution of F1-scores for each model, distinguishing between the datasets used. These graphical results show that the VQC model performs best on the European dataset, with a median F1-score of around 0.75 and moderate variance, which aligns with our results. This suggests that the VQC model can effectively capture the characteristic patterns of real fraudulent transactions. On the BankSim dataset, its performance is slightly lower, with a median of around 0.65 and greater variance, possibly reflecting sensitivity to the synthetic nature of the data.
The SQNN model performs remarkably well and consistently across both datasets, with a median close to 0.8 on BankSim. However, the results on the European dataset show greater variability, which could indicate increased sensitivity to specific architectural combinations. The stability observed on BankSim, in contrast, suggests better generalization. The EQNN model performs significantly worse, with a median F1-score of around 0.4 on both datasets. The low dispersion observed, combined with a few outliers, suggests limited adaptability, which reduces its effectiveness in complex financial fraud detection scenarios.
This statistical analysis reinforces the validity of our comparative study and highlights the critical importance of QML model selection in financial fraud detection. It also underscores the need to systematically explore different combinations of feature maps and ansatz within quantum model architectures. Our conclusions are thus based not only on empirical performance averages but also on statistically significant differences confirmed through rigorous analysis of variance.
Figure 18: Boxplot shows the distribution of the F1-scores obtained by the three QML models evaluated on the BankSim and European datasets. Each box shows the distribution of F1-scores across ten repeated runs for each model.
V-F Noise Analysis
[…]
Figure 19: Impact of quantum noise on the SQNN model on the BankSim dataset.
[…]
Figure 20: Impact of quantum noise on the VQC model on the European dataset.
This comparative analysis underscores the critical role of model architecture in determining quantum noise resilience. While the SQNN model is more susceptible to circuit-level perturbations, the VQC model maintains greater stability and adaptability, making it a more suitable candidate for deployment in real-world quantum applications such as fraud detection.
V-G Computational Complexity of QML Architectures
From a theoretical perspective, the three QML architectures examined in this study, VQC, SQNN, and EQNN, share the same asymptotic computational complexity on quantum hardware. For a parametrized quantum circuit with $n$ qubits and depth $d$, a forward pass requires $O(d\cdot n)$ gates, while a gradient-based update with $p$ trainable parameters and $m$ measurement shots scales as $O(p\cdot m\cdot d\cdot n)$. Under these standard assumptions for variational quantum circuits, all three architectures belong to the same polynomial-time complexity class [8].
Beyond this ideal setting, additional overhead arises from mapping logical circuits to the device’s native gate set and connectivity graph. Gate decomposition, routing, and SWAP insertion may increase circuit depth by a constant or topology-dependent factor, but these hardware effects do not modify the overall asymptotic scaling.
Classical simulation, in contrast, incurs exponential complexity: a state-vector simulator must track $2^{n}$ amplitudes [51, 33], resulting in a cost of $O(2^{n}\cdot d)$ for all three architectures. Practical runtime differences therefore originate not from circuit depth or qubit count, but from how each architecture processes measurements. VQC and SQNN rely on the Sampler primitive to produce probability distributions, whereas EQNN uses the Estimator primitive to compute expectation values. Because each model applies the same feature map and ansatz, their gate-level complexity is identical; only constant-factor differences arise from measurement and output-mapping procedures such as parity mapping in VQC, multi-class probability construction in SQNN, or expectation-value evaluation in EQNN.
Since optimizer convergence depends on the landscape of the loss function, iteration counts are excluded from the asymptotic analysis. Overall, these considerations show that observed runtime differences across architectures arise mainly from measurement primitives, hardware-induced overhead, and classical post-processing rather than from intrinsic differences in computational complexity.
V-H Discussion
The analytical results highlight the strong interaction between feature maps, ansatz configurations, and dataset characteristics in shaping model performance. A model’s ability to balance its performance metrics depends on how effectively the feature map encodes relevant relationships and how well the ansatz can exploit these encoded structures. These interactions vary across datasets and architectures, underscoring the need for configuration-aware model design.
Performance differences across architectures reflect their distinct output-processing mechanisms. VQC consistently benefits from expressive feature map–ansatz combinations, achieving high precision and stable F1-scores, particularly on the European dataset. SQNN demonstrates strong adaptability, excelling with configurations that balance computational simplicity and representational power, such as the Z feature map combined with Real Amplitudes on the BankSim dataset. In contrast, EQNN shows difficulty capturing complex patterns, likely due to structural limitations such as low circuit expressiveness, restricted feature mixing, and the constraints of mapping expectation values to class labels. Its sensitivity to configuration choices further indicates limited architectural flexibility compared to VQC and SQNN. Improving EQNN performance would require increasing circuit depth or expressivity, exploring alternative expectation-to-label mappings, or adopting task-specific loss functions that enhance sensitivity to minority-class patterns. These considerations highlight the importance of aligning architectural properties with appropriate encoding and variational strategies when deploying QML models for fraud detection.
The findings further indicate that entanglement-enhanced feature maps (e.g., ZZ, Pauli) can improve performance, but only when the model can effectively handle the added circuit complexity. SQNN, for instance, benefits substantially from such feature maps when paired with flexible ansatz like TwoLocal. Simpler feature maps, such as Z, frequently achieve strong results due to low depth, reduced noise sensitivity, and easier optimization, making them effective for less expressive models or noisy conditions.
Overall, the differences across feature maps and ansatz configurations align with theoretical principles related to circuit expressivity, entanglement structure, and the coupling between data encoding and variational layers. Configurations that enable richer expressibility and stronger feature mixing yield more stable and accurate models, while shallow or weakly entangling designs limit learning capacity. Although this study is empirical, the trends observed are consistent with established results on expressibility, trainability, and encoding efficiency.
These empirical observations are further supported by statistical evidence from the ANOVA analysis. Model type, feature map, and ansatz all exhibit statistically significant effects on F1-score, with model type producing the highest F-value. This confirms that the observed performance variability arises from meaningful architectural and configuration factors. In contrast, the dataset factor is not statistically significant, indicating that performance trends remain consistent across both the European and BankSim datasets.
To further contextualize these findings, the robustness trends under quantum noise offer additional insight into how architectural choices influence stability in realistic quantum environments. Models with well-structured variational layers demonstrate a stronger ability to preserve learned decision boundaries under perturbations, whereas architectures with limited expressiveness or weaker feature mixing degrade more rapidly. Across the simulated noise channels, depolarizing, phase damping, and amplitude damping exerted the strongest negative effects, while phase-flip noise was comparatively milder. Strategies such as reducing circuit depth, employing noise-adaptive ansatz designs, or integrating error-mitigation techniques may help improve robustness, although potentially at added computational cost.
This study demonstrates the importance of a strategic, data-driven approach to configuring QML models. The interplay between dataset properties, architectural structure, and feature map–ansatz combinations determines the achievable balance between computational efficiency and predictive accuracy. These insights can guide the design of more effective quantum and hybrid quantum models for financial fraud detection and similarly complex decision-making tasks.